OpenStudy (anonymous):

find the no. of ways in which 8 people can be seated at 2 round tables with 4 people sitting at each table?

OpenStudy (anonymous):

got it???

OpenStudy (anonymous):

acually its like ... choose four people ...then arrange them .... permutaion means choosing and aranging

OpenStudy (hba):

36

OpenStudy (anonymous):

that dot means multiplty

OpenStudy (anonymous):

why does the answer in my solution book say this : 7x6x5x3x2x1x2

OpenStudy (anonymous):

wait a sec .... let me take a look at question again

OpenStudy (hba):

WAIT I HAVE A WAY

OpenStudy (anonymous):

@hba having epiphanies are we?

OpenStudy (hba):

IF UR ASKED OF A ROW THEN n! IF UR ASKED ABOUT A CIRCLE OR ROUND ITS (n-1)! AND IF ITS A NECKLACE OR GARLAND ITS 1/2(n-1)!

OpenStudy (hba):

IN UR CASE IT WOULD BE (8-1)! =7!

OpenStudy (hba):

AND 2 TABLES SO U MULTIPLY IT BY 2 THEN

OpenStudy (anonymous):

naaaa

OpenStudy (hba):

Remember that a circular permutation is (n – 1)! Break the seating process into "activities" like this: 1) Choose the people who are seated at each table. C(8,4 = 8•7•6•5/4•3•2•1 = 70). When you choose for 1 table the other four automatically go to the other table. 2) Arrange the 4 people at table A: (4 – 1)! = 3! = 6 3) Arrange the 4 people at table B. Also 3! =6 So answer is C(8,4)•3!•3! = 70•6•6 = 2520

OpenStudy (anonymous):

he copied it from somewhere :D

OpenStudy (anonymous):

anyway ... nice work @hba :D

OpenStudy (hba):

You are free: to Share — to copy, distribute and transmit the work to Remix — to adapt the work

OpenStudy (anonymous):

^^ LOL

OpenStudy (anonymous):

hey @hba so you used combinations and not only permutations for this questions?

OpenStudy (hba):

yeah